A non-uniform bar rests across two supports that are 20 ft. apart. When loads of 200 lb, 4 ft from

     end  A and 100 lb, 6 ft  from end B, are on the bar, end  A supports 233 lb and end B supports 197 lb.

     Find the weight of the bar and the position of the center of gravity.


1
Expert's answer
2021-05-09T13:05:04-0400
"W=233+197-200-100=130\ lb\\
200(4)+130x+100(20-6)=233(0)+197(20)\\x=13.4\ ft \text{ from A}"